[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Oct 15, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Oct 15 20:55:49 UTC 2017


HSU Shift Log for Sun Oct 15 2017 16:55:44 GMT-0400 (EDT)

Shift: H

Forecaster: D Zelinsky

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Microwave imagery stopped ingesting sometime last night  
(ZGOES-R)

AWIPS-2/NCP issues: None

Other Techincal issues: None

Miscellaneous issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171015/3a546b8a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list